Suppose that the length of a rectangle is 3 inches longer than the width and that the perimeter of the rectangle is 78.
a) set up an equation involving only w, the width of the rectangle.
b) solve this equation algebraically to find the length of the rectangle. Find the width as well. Use the appropriate units of measure in your answers.
Length_____, Width______

Answer by checkley77(12844)

L=W+3
2L+2W=78
2(W+3)+2W=78
2W+6+2W=78
4W=78-6
4W=72
W=72/4
W=18 ANS. FOR THE WIDTH.
L=18+3=21 ANS. FOR THE LENGTH.
pROOF:
2*21+2*18=78
42+36=78
78=78